# Discovering Large-Scale Structure at in the C3VO Survey

### v1

@@ 4.5 Finding a Mass Relation
The protostructure candidates in the VMC maps, generated from both the real and mock data, have various measures of galaxy overdensity. These measures result from the Gaussian fit to galaxy overdensity as a function of redshift, or, equivalently, VMC slice, for a given candidate associated with a given structure (see §3.3). The Gaussian fit also can give us a sense of the candidate detection strength through parameters such as the integrated area or amplitude of the fit. With the total masses provided through the lightcone data and the candidates we find in the mock VMC maps, we can attempt to use one or more of the candidate detection strength parameters as a proxy for the structure total mass.
However, some structures in the lightcone are close enough together such that we cannot distinguish them as separate structures in the mock VMC maps, and so the candidates we detect will be some mixture of the signal from multiple structures. In such cases, the candidate detection strength will not be representative of the “known” structure mass, resulting in additional scatter in the relation between structure mass and its proxy. In order to mitigate the effect of these ambiguous cases, we impose an isolation criterion on the subsumed structures for the exercise described in this section. Starting from the highest mass subsumed structure, we check for any other structure within its effective radius. The effective radius is defined as the average radius that encloses 90% of the stellar mass of a structure and is dependent on mass of the subsumed structure, as well as redshift. We estimate the effective radius using the relations given in Figure 2 of Muldrew et al. (2015), where the effective radii are given as a function of redshift in three different mass bins of , , and . For a given redshift, the range of effective radii can be considerably large (up to 10 Mpc in comoving space for the largest mass bin), but for simplicity we adopt the mean value reported. Note that, for this exercise, we assume that the total mass of the structure as reported in the GAEA lightcone is approximately equal to the value used by Muldrew et al. (2015).
@@
where is the transverse separation, is the redshift separation, is the scaling factor, and is the effective radius. For the masses and redshifts of the structures in the lightcone, the effective radii typically fall somewhere between 2 to 6 Mpc. It is worth noting that imposing this isolation criterion is a small effect, removing only about 1% of structures between in the lightcone. If we do not find any other structures within this ellipsoid with masses greater than 10% of the seed structure, then we consider that structure to be isolated. Of the subsumed structures in the lightcone, the vast majority (99%) are considered to be isolated by this metric.
@@
For each isolated subsumed structure that has a corresponding detection in the VMC maps of the mock galaxies, we compared both the integrated area of the Gaussian fit to the candidate and the amplitude of the same fit to the mass of that structure for each SF variant in the mocks. Of the two potential proxies, we find stronger correlation between the mass and the Gaussian amplitude and thus adopt it as our generalized mass proxy. More specifically, we perform a Spearman correlation test on of isolated structure versus the Gaussian amplitudes of the associated candidates, finding correlation coefficients of for the SF0.5, SF1.0 IRAC124, SF1.0, SF1.5, SF2.0, and SF4.0 mocks777For those SF variants where multiple mocks were created, the structures and candidates across all mocks of a given variant were combined together prior to running a Spearman test. This is also the case for the fits described below., respectively. The Spearman test returned a rejection of the null hypothesis of no correlation at the 3 level for every SF variants except SF0.5, where the rejection of the null hypothesis was at the 3.5 level. In Figure 9 we show the mass versus the Gaussian amplitude of candidates associated with isolated structures in the mocks for each of the SF variants.
@@
We attempt to construct a mass relation by fitting the Gaussian amplitudes, , of our candidates associated with structures in the mocks with their known masses in logarithmic space to the functional form , where and . The fit of Gaussian amplitude () versus is performed for each SF variant, with dramatic differences in both the fit parameters and the scatter around the relation for the different variants (see Table 2) for the resultant parameters. The large scatter in the data in all of the SF variants tested here results in lower mass structures having their masses overestimated by these relations and higher mass structures having their masses underestimated (Fig. 9). While this issue is the least acute in the SF4.0 case, the fit still precludes the possibility of returning masses in excess of 1014.6 , despite 10 structures in this region of the lightcone having masses that exceed that threshold (Fig. 10). Even in the SF4.0 case, the scatter is considerable, with mass estimates being no less certain than 0.3 dex.
Mock | ||||
SF0.5 | 247.1 | -1448.8 | 7627.1 | -3.3 |
SF1.0aaCut at IRAC124, all other mocks are cut at IRAC124.8 | 14.7 | -1398.3 | 1400.2 | -1080.6 |
SF1.0 | 15.0 | -5943.5 | 5945.6 | -2991.4 |
SF1.5 | 234.8 | -1266.4 | 7145.8 | -3.1 |
SF2.0 | 14.9 | -4349.8 | 4351.9 | -2545.1 |
SF4.0 | 14.7 | -8432.3 | 8434.4 | -7193.2 |
@@
Although we find some utility in performing this mass fit, the estimated masses resulting from the fit are of limited use as they lack both accuracy and precision, especially at the lower SF values. As such, we do not report masses in our final protostructure candidate catalog but simply provide the Gaussian amplitude of the candidates in our catalog as well as in most of the figures in this paper. The conversions in Table 2 are provided for readers wishing to assign values to our candidates, with the understanding that the resultant values are likely biased at some level and lacking in precision. Zoom-in three-dimensional mapping of each structure, like those performed in Cucciati et al. (2018); Forrest et al. (2023); Shah et al. (2024a); Staab et al. (2024), would be required in order to recover more robust mass estimates.
@@
Figure 11 shows the distribution of Gaussian amplitudes and redshifts for our final set of candidates in all three C3VO fields as compared to candidates in the mocks appropriate for each field. For the CFHTLS-D1 field, we combine all combinations of the SF1.0 mocks cut at IRAC124 and the SF0.5 mock. For the COSMOS field, we use all combinations of the SF1.0 mocks cut at IRAC124.8. And for the ECDFS field, we use a combination of the SF2.0 and SF4.0 candidate catalogs as the ECDFS spectroscopic redshift fraction is roughly halfway in between that of the two mocks. Overall, a high degree of concordance is seen between the distributions of the candidates in each field versus those of the appropriate mocks indicating that likely (a) the large-scale structure (LSS) in the lightcone is largely similar to the LSS in the observed fields, and (b) our mapping methodology is equally effective at detecting structure in the mocks as in the real data.
@@
Of course, there are several degeneracies affecting this argument on various levels. Chief among them is that we cannot fully rule out the possibility that the properties of LSS in the observed universe are considerably different than those in our simulations. Such differences could result in discrepancies in the purity and completeness between the real and mock data that then, by chance, produce similar distributions. However, given all the tests performed on the consistency between simulated galaxies and the observed galaxy catalogs and the care taken to reproduce our spectroscopic and photometric data quality and analysis strategy in the mocks, we take this as the less likely scenario and consider the high degree of concordance between the mock candidates and the candidates in the real data as lending credence to the efficacy of our approach.
@@ 4.6 Completeness and Purity
where and are the number of recovered and total subsumed structures respectively. For this calculation, we limit the search area to the central 0.91 by 0.91 deg2 of each mock field to avoid complications with edge effects. In order to further reduce ambiguities, we limit our completeness calculations to only those structures that are isolated.
@@
There is a noticeable drop-off in completeness between the and redshift bins, approximately 30% on average, though the completeness still generally remains high for the most massive structures and most heavily spectroscopically-sampled mocks. A much larger drop in completeness is observed in our highest redshift bin at , where we recover, at most, 20% completeness at (SF4.0). The completeness in the IRAC1 mocks are roughly equivalent to the SF0.5 IRAC1 mock for , but the completeness numbers for the former mock are noticeably worse at and fall to zero at . The latter is a natural consequence of the shallower photometric cut in that mask given how few galaxies are brighter than this apparent magnitude limit at such redshifts.
For the purity analysis, we impose no isolation criterion as we are mainly concerned with recovering signal no matter what form it takes. The purity is given simply
where and are the number of matched and total candidates respectively.
In Figure 13 we plot the recovered purity as a function of SF and as a function of Gaussian amplitude of the candidates. The purity is seen to improve precipitously with spectroscopic fraction up until it levels off at SF1.5. Higher levels of spectroscopic sampling do not appear to improve the purity beyond about 80%. However, the protostructure catalog that is used to assess the purity is cut at , and it is likely that we are detecting lower mass structures at higher SF levels and spuriously considering them as false positives. While there does appear to be a dip in purity between SF1.5 and SF2.0, we note that we only analyzed multiple mocks for the SF1.0 cases. The scatter seen in the purity values across the different mocks for the SF1.0 case, represented by the error bars in the shaded region of Figure 13, would likely be the similar for higher SF variants. Thus, it is likely that the purity numbers for SF1.5, 2.0, and 4.0 are all statistically consistent. While our purity numbers noticeably improve when restricted to higher amplitude candidates (i.e., generally higher mass structures), we impose no cut on our candidate catalog in order to maximize the number of structures we can potentially find.
@@ 5.1 The C3VO Protostructure Candidate Catalog
We provide on the Centre de Données astronomiques de Strasbourg (CDS) website888https://cds.u-strasbg.fr/ an electronic table of all 565 unique protostructure candidates in the galaxy density maps of the COSMOS, CFHTLS-D1, and ECDFS fields that fulfill all criteria outlined in §3.3 over the redshift range 999Due to decreased data quality, in particular the shallower IRAC data in the field, the candidate catalog for the CFHTLS-D1 field is truncated to the redshift range .. Judging from our comparisons with simulations, the protostructure candidates listed likely have masses in excess of , i.e., the progenitors of modern-day groups and clusters.
@@
The candidates in this catalog vary tremendously in their strength, with Gaussian amplitudes spanning over two orders of magnitude from 20 to 2000, which likely indicates a similar dynamic range in protostructure masses, i.e., from to . Of the 565 protostructure candidates, 343 are in the COSMOS field, 142 are in CFHTLS-D1, and 80 are in ECDFS. This distribution is reasonable given the size of each field and the relative data quality. In the CDS table we provide the field associated with each candidate, the barycenter in the RA, Dec, and redshift dimensions, the Gaussian amplitude of the associated candidate, the uncertainty in the Gaussian amplitude, the type of candidate (i.e., single vs. in a multi-peak chain, see §3.3), and a unique identifying designation. Though we urge caution, the Gaussian amplitudes can be translated into masses by adopting the appropriate parameters from Table 2 and the formula presented in §4.5 for each field (see §4.5 for the appropriate mocks for each field). However, better methods exist for estimating the mass of candidate protostructures through dedicated zoom-in mapping (see, e.g., Cucciati et al. 2018; Lemaux et al. 2018; Shen et al. 2021; Forrest et al. 2023; Shah et al. 2024a; Staab et al. 2024).

### v2

@@ 4.5 Finding a Mass Relation
The protostructure candidates in the VMC maps, generated from both the real and mock data, have various measures of galaxy overdensity. These measures result from the Gaussian fit to galaxy overdensity as a function of redshift, or, equivalently, VMC slice, for a given candidate associated with a given structure (see Section 3.3). The Gaussian fit also can give us a sense of the candidate detection strength through parameters such as the integrated area or amplitude of the fit. With the total masses provided through the lightcone data and the candidates we find in the mock VMC maps, we can attempt to use one or more of the candidate detection strength parameters as a proxy for the structure total mass.
However, some structures in the lightcone are close enough together such that we cannot distinguish them as separate structures in the mock VMC maps, and so the candidates we detect will be some mixture of the signal from multiple structures. In such cases, the candidate detection strength will not be representative of the “known” structure mass, resulting in additional scatter in the relation between structure mass and its proxy. In order to mitigate the effect of these ambiguous cases, we impose an isolation criterion on the subsumed structures for the exercise described in this Section. Starting from the highest mass subsumed structure, we check for any other structure within its effective radius. The effective radius is defined as the average radius that encloses 90% of the stellar mass of a structure and is dependent on mass of the subsumed structure, as well as redshift. We estimate the effective radius using the relations given in Figure 2 of Muldrew et al. (2015), where the effective radii are given as a function of redshift in three different mass bins of , , and . For a given redshift, the range of effective radii can be considerably large (up to 10 Mpc in comoving space for the largest mass bin), but for simplicity, we adopt the mean value reported. Note that, for this exercise, we assume that the total mass of the structure as reported in the GAEA lightcone is approximately equal to the value used by Muldrew et al. (2015).
@@
where is the transverse separation, is the redshift separation, is the scaling factor, and is the effective radius. For the masses and redshifts of the structures in the lightcone, the effective radii typically fall somewhere between 2 and 6 Mpc. It is worth noting that imposing this isolation criterion is a small effect, removing only about 1% of structures between in the lightcone. If we do not find any other structures within this ellipsoid with masses 10% of the seed structure, then we consider that structure to be isolated. Of the subsumed structures in the lightcone, the vast majority (99%) are considered to be isolated by this metric.
@@
For each isolated subsumed structure that has a corresponding detection in the VMC maps of the mock galaxies, we compared both the integrated area of the Gaussian fit to the candidate and the amplitude of the same fit to the mass of that structure for each SF variant in the mocks. Of the two potential proxies, we find stronger correlation between the mass and the Gaussian amplitude and thus adopt it as our generalized mass proxy. More specifically, we perform a Spearman correlation test on of isolated structure versus the Gaussian amplitudes of the associated candidates, finding correlation coefficients of for the SF0.5, SF1.0 IRAC124, SF1.0, SF1.5, SF2.0, and SF4.0 mocks777For those SF variants where multiple mocks were created, the structures and candidates across all mocks of a given variant were combined together prior to running a Spearman test. This is also the case for the fits described below., respectively. The Spearman test returned a rejection of the null hypothesis of no correlation at the 3 level for every SF variant except SF0.5, where the rejection of the null hypothesis was at the 3.5 level. In Figure 9 we show the mass versus the Gaussian amplitude of candidates associated with isolated structures in the mocks for each of the SF variants.
@@
We attempt to construct a mass relation by fitting the Gaussian amplitudes, , of our candidates associated with structures in the mocks with their known masses in logarithmic space to the functional form , where and which was previously used in Hung et al. (2020). The fit of Gaussian amplitude () versus is performed for each SF variant, with dramatic differences in both the fit parameters and the scatter around the relation for the different variants (see Table 2) for the resultant parameters. The large scatter in the data in all of the SF variants tested here (0.3 dex) results in lower-mass structures having their masses overestimated by these relations and higher-mass structures having their masses underestimated (Fig. 9). While this issue is the least acute in the SF4.0 case, the fit still precludes the possibility of returning masses in excess of 1014.6 , despite 10 structures in this region of the lightcone having masses that exceed that threshold (Fig. 10). Even in the SF4.0 case, the scatter is considerable, with mass estimates being no less certain than 0.3 dex.
Mock | aaScatter given in units of , as defined in Lemaux et al. (2022) and references therein. | ||||
SF0.5 | 141.9 | -995.9 | 4655.9 | -3.2 | 0.32 |
SF1.0bbCut at IRAC124, all other mocks are cut at IRAC124.8 | 14.5 | -2867.4 | 2869.1 | -2867.0 | 0.35 |
SF1.0 | 14.3 | -10313.5 | 10314.7 | -15714.5 | 0.36 |
SF1.5 | 235.3 | -1286.6 | 7260.7 | -3.1 | 0.34 |
SF2.0 | 14.8 | -9761.1 | 9763.1 | -6476.6 | 0.31 |
SF4.0 | 14.8 | -4431.6 | 4433.72 | -3731.6 | 0.32 |
@@
Although we find some utility in performing this mass fit, the estimated masses resulting from the fit are of limited use as they lack both accuracy and precision, especially at the lower SF values. As such, we do not report masses in our final protostructure candidate catalog but simply provide the Gaussian amplitude of the candidates in our catalog as well as in most of the Figures in this paper. The conversions in Table 2 are provided for readers wishing to assign values to our candidates, with the understanding that the resultant values are likely biased at some level and lacking in precision. Zoomed-in three-dimensional mapping of each structure, like those performed in Cucciati et al. (2018), Forrest et al. (2023), Shah et al. (2024a), and Staab et al. (2024), would be required in order to recover more robust mass estimates.
@@
Figure 11 shows the distribution of Gaussian amplitudes and redshifts for our final set of candidates in all three C3VO fields as compared to candidates in the mocks appropriate for each field. For the CFHTLS-D1 field, we combine all combinations of the SF1.0 mocks cut at IRAC124 and the SF0.5 mock. For the COSMOS field, we use all combinations of the SF1.0 mocks cut at IRAC124.8. And for the ECDFS field, we use a combination of the SF2.0 and SF4.0 candidate catalogs as the ECDFS spectroscopic redshift fraction is roughly halfway in between that of the two mocks. Overall, a high degree of concordance is seen between the distributions of the candidates in each field versus those of the appropriate mocks indicating that likely (a) the LSS in the lightcone is largely similar to the LSS in the observed fields, and (b) our mapping methodology is equally effective at detecting structure in the mocks as in the real data.
@@
Of course, there are several degeneracies affecting this argument on various levels. Chief among them is that we cannot fully rule out the possibility that the properties of LSS in the observed Universe are considerably different than those in our simulations. Such differences could result in discrepancies in the purity and completeness between the real and mock data that then, by chance, produce similar distributions. However, given all of the tests performed on the consistency between simulated galaxies and the observed galaxy catalogs and the care taken to reproduce our spectroscopic and photometric data quality and analysis strategy in the mocks, we take this as the less likely scenario and consider the high degree of concordance between the mock candidates and the candidates in the real data as lending credence to the efficacy of our approach.
@@ 4.6 Completeness and Purity
where and are the number of recovered and total subsumed structures, respectively. For this calculation, we limit the search area to the central 0.91 0.91 deg2 of each mock field to avoid complications with edge effects. In order to further reduce ambiguities, we limit our completeness calculations to only those structures that are isolated.
@@
There is a noticeable drop-off in completeness between the and redshift bins, approximately 30% on average, though the completeness still generally remains high for the most massive structures and most heavily spectroscopically sampled mocks. A much larger drop in completeness is observed in our highest redshift bin at , where we recover, at most, 20% completeness at (SF4.0). The completeness in the IRAC1 mocks are roughly equivalent to the SF0.5 IRAC1 mock for , but the completeness numbers for the former mock are noticeably worse at and fall to zero at . The latter is a natural consequence of the shallower photometric cut in that mask given how few galaxies are brighter than this apparent magnitude limit at such redshifts.
For the purity analysis, we impose no isolation criterion, as we are mainly concerned with recovering signal no matter what form it takes. The purity is given simply as
where and are the number of matched and total candidates, respectively.
In Figure 13 we plot the recovered purity as a function of SF and as a function of Gaussian amplitude of the candidates. The purity is seen to improve precipitously with spectroscopic fraction up until it levels off at SF1.5. Higher levels of spectroscopic sampling do not appear to improve the purity beyond about 80%. However, the protostructure catalog that is used to assess the purity is cut at , and it is likely that we are detecting lower-mass structures at higher SF levels and spuriously considering them as false positives. While there does appear to be a dip in purity between SF1.5 and SF2.0, we note that we only analyzed multiple mocks for the SF1.0 cases. The scatter seen in the purity values across the different mocks for the SF1.0 case, represented by the error bars in the shaded region of Figure 13, would likely be the similar for higher SF variants. Thus, it is likely that the purity numbers for SF1.5, 2.0, and 4.0 are all statistically consistent. While our purity numbers noticeably improve when restricted to higher-amplitude candidates (i.e., generally higher mass structures), we impose no cut on our candidate catalog in order to maximize the number of structures we can potentially find.
@@ 5.1 The C3VO Protostructure Candidate Catalog
We provide in Table 3 an abbreviated list of all 561 unique protostructure candidates in the galaxy density maps of the COSMOS, CFHTLS-D1, and ECDFS fields that fulfill all criteria outlined in Section 3.3 over the redshift range 888Due to decreased data quality, in particular the shallower IRAC data in the field, the candidate catalog for the CFHTLS-D1 field is truncated to the redshift range .. The abbreviated 623 individual peaks and their parameters are reported in Table 4. The full tables can be found online. Judging from our comparisons with simulations, the protostructure candidates listed likely have masses in excess of , i.e., the progenitors of modern-day groups and clusters.
Field | Candidate IDaa“MP” and “SP” refer to Gaussian fits originating from multiple and single peak chains, respectively. The Gaussian amplitudes are less reliable in the case of multiple peak chains due to the blended flux values between neighboring peaks (see Fig. 3). | PeaksbbNumber of peaks subsumed together within a transverse radius of 2 Mpc and . | R.A. | Decl. | Amp.ccAmplitude of the Gaussian fit to the linked candidate. If the candidate consists of multiple peaks, the reported amplitude is the summed total of the amplitudes of the constituent peaks. | |
COSMOS | C_SP1 | 1 | 150.1993 | 2.3781 | 2.462 | 2013 32 |
COSMOS | C_SP2 | 1 | 150.0125 | 2.1800 | 2.432 | 1640 18 |
COSMOS | C_SP3 | 1 | 150.3551 | 2.3534 | 4.573 | 1179 14 |
COSMOS | C_MP5 | 1 | 150.3451 | 2.6811 | 2.750 | 559 47 |
CFHTLS-D1 | V_MP1 | 1 | 36.3824 | -4.2238 | 3.754 | 2183 427 |
CFHTLS-D1 | V_MP2 | 1 | 36.4027 | -4.2737 | 3.884 | 1739 84 |
CFHTLS-D1 | V_SP2 | 1 | 36.6969 | -4.3563 | 3.292 | 1574 27 |
ECDFS | E_SP1 | 1 | 53.1419 | -27.8795 | 3.469 | 1731 58 |
ECDFS | E_SP2 | 1 | 53.1926 | -27.9161 | 3.080 | 1304 18 |
ECDFS | E_SP3 | 1 | 53.0948 | -27.8190 | 2.813 | 1279 17 |
Note. — Candidates within 2 Mpc of the boundary of our search region are denoted by the symbol.
Field | Peak IDaaPeaks from multiple peak chains are denoted by a superscript, which are numbered by the relative detection strength of the peak in the chain ordered from high to low. | Candidate IDbbName of the protostructure candidate that the peak was assigned. The naming conventions of candidates follow those of Table 3. | R.A. | Decl. | ccRedshift dispersion of the Gaussian fit to the linked candidate. Uncertainties are the formal random error derived from the covariance matrix of the fit. Value is less certain in cases where peak is embedded in a multiple peak chain. | Amp.ddfootnotemark: | |
COSMOS | C_SP182 | 150.3024 | 2.1370 | 2.012 0.000 | 0.003 0.000 | 93 7 | |
COSMOS | C_SP62 | 149.9838 | 2.0783 | 2.012 0.000 | 0.005 0.000 | 192 7 | |
COSMOS | C_SP213 | 149.9300 | 1.9182 | 2.018 0.001 | 0.010 0.002 | 71 4 | |
COSMOS | C_SP38 | 150.4395 | 2.3726 | 2.024 0.000 | 0.010 0.000 | 265 6 | |
CFHTLS-D1 | V_SP67 | 36.1684 | -4.8729 | 2.008 0.002 | 0.004 0.001 | 90 9 | |
CFHTLS-D1 | V_SP84 | 36.1129 | -4.4638 | 2.038 0.001 | 0.007 0.001 | 63 4 | |
CFHTLS-D1 | V_SP106 | 36.6670 | -4.2870 | 2.046 0.001 | 0.007 0.002 | 48 4 | |
ECDFS | E_SP21 | 53.1293 | -27.8211 | 2.005 0.010 | 0.010 0.008 | 168 47 | |
ECDFS | E_SP71 | 52.9692 | -27.8606 | 2.014 0.011 | 0.014 0.013 | 75 10 | |
ECDFS | E_SP48 | 53.0669 | -27.7158 | 2.025 0.009 | 0.011 0.010 | 110 15 |
Note. — Peaks and candidates within 2 Mpc of the boundary of our search region are denoted by the symbol.
@@
The candidates in this catalog vary tremendously in their strength, with Gaussian amplitudes spanning over 2 orders of magnitude from 20 to 2000, which likely indicates a similar dynamic range in protostructure masses, i.e., from to . Of the 561 protostructure candidates, 343 are in the COSMOS field, 138 are in CFHTLS-D1, and 80 are in ECDFS. This distribution is reasonable given the size of each field and the relative data quality. In Table 3 we provide the field associated with each candidate, the barycenter in the R.A., decl., and redshift dimensions, the Gaussian amplitude of the associated candidate, the uncertainty in the Gaussian amplitude, the number of subsumed peaks and type of candidate (i.e., single versus in a multipeak chain, see Section 3.3), and a unique identifying designation. Though we urge caution, the Gaussian amplitudes can be translated into masses by adopting the appropriate parameters from Table 2 and the formula presented in Section 4.5 for each field (see Section 4.5 for the appropriate mocks for each field). Better methods exist for estimating the mass of candidate protostructures through dedicated zoom-in mapping (see, e.g., Cucciati et al. 2018; Lemaux et al. 2018; Shen et al. 2021; Forrest et al. 2023; Shah et al. 2024a; Staab et al. 2024). However, these methods are highly parameter-sensitive and require each structure to be examined on an individual basis, with careful consideration for various factors like the spectral sampling, the tracer population, and the galaxy bias. Thus, they cannot be broadly applied to a large sample such as ours.